Shift summary, Fri 24 Dec 99

Arun Tripathi (arun@physics.ucla.edu)
Fri, 24 Dec 1999 05:02:33 -0800 (PST)



Shift Summary for owl shift, 24 Dec 99

On Shift: Y. B. Hsiung, Arun T.

Hours of beam available at all intensities: 40 min
Hours of beam available at nominal intensity: 40 min
Hours of E799 data (special runs not included): 40 min
Hours of special runs (laser scans, muon runs...): 1 hr (Laser Scan)

Accounting of lost time:

========================================================================
Runs ended this shift:
Run 15279 - 9,25M events - BEAM_799
Run 15280 - Laser Scan

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: E799 datataking until 00:43, and then beam goes away
due to a quench and a bad Kautzky valve in the F1 sector of the TeV.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:43 TeV quench at F1, est. 15 minutes.
15279 1:10 TeV cryo problems at F1, est. 1 hr. Stop BEAM_799
run 15279 to utilize the no beam time for a
laser scan.
15280 1:20 Start the laser scan run 15280.
2:01 Laser scan finished.
2:05 Start BEAM_799 run in anticipation of beam, but
no beam yet.
2:45 Stuck Kautzky valve at F1. Experts en-route, est.
3 hrs.
4:55 Kautzky valve is fixed. Cooling down F1, est. 1.5
hrs
6:30 Laser illumination plot shows missing CsI top-left
quadrant, reloaded ET, fixed.
7:00 No beam yet.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift